(–a) + b = b + Additive inverse of __________. 
 (–a) + b = b + Additive inverse of __________. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

a:

(–a) + b = b + (-a)
(–a) + b = b + Additive inverse of (a)

Thanks